Child pages
  • 202103 - Public report
Skip to end of metadata
Go to start of metadata


Unscheduled:

  • March 26th: There was a problem with login to edusign due to faulty metadata which now has been fixed.
  • March 18th: The link between SUNET (STHB) and KI was down. CFP restart from PFE on sthb-r1 router helped to restore the link.
  • March 10th: The link between Trollhättan and Göteborg was down after the fiber work from SUNETTICKET-7012. Fiber provider has fixed the fiber and the circuits was restored.
  • March 8th: Dark fiber between Östersund (OSD) and MIUN was down due to a fiber cut.
  • March 6th: Link was down towards new rise connection in Jönköping. The connection is restored.
  • March 1st: The router hh-r1.sunet.se was unreachable. Two router components were replaced and the router restarted. The equipment is up and running for almost a week.
  • February 2nd: The router slu-r2 was unreachable for approximately 10 minutes. Reason for outage not found.
  • January 24th: Power outage at the site initially caused router to go down. After power was back, FPC did not work as expected and had to be replaced. FPC was replaced and the reachability to the router was restored.

Scheduled:

  • March 31st: CFP et-9/1/1 in tug-r1 was replaced in order to comply with the original setup.
  • March 31st: Supplier performed fiber measurement and cleaned connectors to improve fiber values on the fiber between bos-c2 and bos-r2
  • March 25th: Provider performed planned fiber maintenance in Junosuando.
  • March 25th: Supplier performed fiber maintenance in the stretch between Hjälta -Norrtjärn and Norrtjärn - Tuggen.
  • March 24th: Tele2 perform rerouting maintenance. This affected the fiber stretch LND88 (Lund) and LYB (Lyby).
  • March 23rd: We needed to perform maintenance on lnu-r2 in order to fix a minor alarm on the router.
  • March 23rd: Supplier performed fiber measurement and clean connectors to improve fiber values on the fiber between bos-c2 and bos-r2 affecting only Rise.
  • March 22nd: Provider performed planned relocation of fiber cable maintenance.
  • March 22nd: Supplier performed fiber measurement and clean connectors to improve fiber values on the fiber between bos-c2 and bos-r2 affecting only Rise.
  • March 10th: Supplier performed fiber measurement and clean connectors to improve fiber values on the fiber between bos-c2 and bos-r2.
  • March 9th: The fiber between GBG7 (Göteborg) - TRH (Trollhättan) was to be relocated. New stretch GBG7 (Göteborg) - BHCK (Backa) outages were expected. The service is down beyond the end of maintenance window.
  • March 8th: Provider performed software upgrade on a core router in their network. This affected the VPN-NNI with Tele2 in Göteborg.
  • March 2nd: Provider performed a mandatory maintenance, reworking circuits which affected the link between LBA (Örebro)- BBG (Björneborg)

Scope

Unscheduled

Scheduled

Total

Hardware

2 pcs 2d 18:30

0 pcs

2 pcs 2d 18:30

Link

3 pcs 3d 14:34

11 pcs 14:56

14 pcs 4d 04:22

None

2 pcs 01:02

2 pcs 07:44

4 pcs 08:46

Routing

0 pcs

0 pcs

0 pcs

Software

1 pcs 1:54

0 pcs

1 pcs 1:54

  • No labels